Linear System
Solve each linear system using the method of substitution.
a) y=6-3x
y=2x+1
b) 5x-y=4
3x+y=4

Answer by MathTherapy(10858)

Y = 6 - 3x
Y = 2x + 1
Substitute the y-value in eq (i) into eq (ii). This gives us:
6 - 3x = 2x + 1
- 5x = - 5
x = 1
Substitute this x-value in eq (ii), we get: y = 2(1) + 1
Y = 3
Answer – (1, 3)
You should now be able to do the next one. Good luck!!